# BlueSuffix: Reinforced Blue Teaming for Vision-Language Models Against Jailbreak Attacks

### v1

@@ Appendix C LLM-based Rewrite Template
Appendix D Suffix Generator
@@
Figures 7 and 8 demonstrate that our suffix generator produces diverse outputs while maintaining a high degree of compatibility with benign prompts. In Figure 7, the attacker inputs image-text prompts. Our BlueSuffix purifies the visual prompts and rewrites the textual prompts by appending a suffix. With the purified prompt, the target VLM successfully identifies the malicious content and generates a benign response, including an explanation for refusing to answer. Figure 8 shows a user asking the target VLM benign queries. Our BlueSuffix processes these queries effectively, particularly for the textual prompts. The suffix generator produces a positive response that is relevant to the textual prompt, guiding the target VLM in answering the question. Our method focuses on enabling VLMs to independently evaluate queries, making it more practical for real-world scenarios.
@@
Appendix E Additional Results of the Ablation Studies
Here, we present the detailed results of our ablation studies, shown in Tables 6 and 7. Additionally, Figure 9 illustrates an example of a strong jailbreak textual prompt rewritten by GPT-4o.
Table 6: The detailed ASR (%) results of our ablation studies on the open-source VLMs.
Jailbreak Topics Target VLMs LLaVA-v1.5-7B MiniGPT-4 Text Purifier Text Purifier + Image Purifier + Image Purifier + Text Purifier Text Purifier + Image Purifier + Image Purifier + Suffix Generator Text Purifier Suffix Generator Suffix Generator Text Purifier Suffix Generator IA 24.74 17.53 25.77 17.53 22.68 17.53 26.80 24.74 HS 15.34 15.95 16.56 9.82 20.68 17.79 22.09 13.50 MG 50.00 18.18 47.73 50.00 27.27 25.00 18.18 20.45 PH 31.94 22.92 29.17 52.78 19.44 15.97 22.92 26.39 EH 20.49 13.11 11.48 13.11 16.39 8.20 12.30 10.66 FR 35.71 20.13 31.82 21.43 30.52 23.38 24.68 24.68 PO 13.76 11.01 7.34 14.68 10.09 11.93 11.01 17.43 PL 27.45 15.69 20.26 4.58 16.99 13.07 10.46 4.58 PV 20.86 17.27 18.71 54.68 17.99 15.83 17.99 23.74 LO 13.08 12.31 9.23 6.15 15.38 7.69 6.15 3.85 FA 19.76 12.57 11.38 2.40 20.96 10.18 8.38 1.80 HC 11.93 11.01 0.92 1.83 11.93 10.09 9.17 9.17 GD 6.71 12.75 6.04 8.05 6.71 6.71 4.03 4.70 Average 22.44 15.42 18.19 19.77 18.23 14.11 14.94 14.28
Table 7: The detailed ASR (%) results of our ablation studies on the commercial VLM.
Jailbreak Topics Target VLMs Gemini (LLaVA) Gemini (MiniGPT-4) Text Purifier Text Purifier + Image Purifier + Image Purifier + Text Purifier Text Purifier + Image Purifier + Image Purifier + Suffix Generator Text Purifier Suffix Generator Suffix Generator Text Purifier Suffix Generator IA 2.06 2.06 4.12 2.06 3.12 0.00 2.08 1.03 HS 4.91 4.29 2.45 2.45 2.45 3.68 6.13 3.07 MG 2.27 0.00 2.27 2.27 0.00 2.27 2.27 4.55 PH 1.39 2.08 2.08 2.08 2.08 1.39 2.08 1.39 EH 0.82 0.82 1.64 2.46 0.82 0.00 1.64 0.82 FR 5.84 1.30 1.95 2.60 1.30 1.30 4.55 1.95 PO 2.75 2.75 1.83 3.67 3.67 3.67 1.83 1.83 PL 1.31 0.65 1.96 0.00 0.65 0.00 0.65 0.00 PV 3.60 3.60 5.76 1.44 3.60 2.88 4.32 0.72 LO 0.00 0.00 0.00 0.00 0.77 0.00 0.00 0.00 FA 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 HC 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 GD 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Average 1.92 1.35 1.85 1.46 1.42 1.17 1.97 1.18

### v2

@@ Appendix A LLaMA as the Text Purifier
Appendix B Semantic Similarity between ‘Original prompt’ and ‘Text Purifier + Suffix Generator prompt’
We conduct experiments to assess whether the ‘Text Purifier + Suffix Generator’ alters the original meaning of the instructions. We employed GPT-4o to evaluate the semantic similarity between the original prompts and the prompts processed by our ‘Text Purifier + Suffix Generator’. Specifically, we asked GPT-4o to assign a similarity score on a scale from 0 to 5, where higher scores indicate greater semantic similarity. The evaluation was guided by the following prompt template:
“Given the following Sentence A: ‘{}’ and Sentence B: ‘{}’, your task is to evaluate if these two sentences contain similar semantic elements or share a partially overlapping meaning. Focus on finding any significant overlap in ideas, context, or main topics.”
The scoring criteria were defined as follows:
- **5**: Major idea or context is similar or shared.
- **4**: Somewhat similar meaning with minor phrasing differences.
- **3**: General topic overlap but differing in focus or detail.
- **2**: Some relation but with noticeable differences in meaning.
- **1**: Minimal connection with faint thematic overlap.
- **0**: Completely unrelated.
We performed this semantic comparison on all textual prompts in the MM-SafetyBench dataset and conducted experiments using both GPT-4o and LLaMA as the text purifier. The results are as follows:
- When using GPT-4o as the text purifier, the average similarity score was 4.82.
- When using LLaMA as the text purifier, the average similarity score was 4.87.
These high scores indicate a strong semantic consistency between the original prompts and the processed prompts. Therefore, our ‘Text Purifier + Suffix Generator’ effectively preserves the original meaning of the instructions while enhancing the model’s robustness against jailbreak attacks.
@@ Appendix C LLM-based Rewrite Template
Appendix D LLM-based Judge Prompt
Figure 6 illustrates the LLM-based judge prompt.
Figure 6: The LLM-based judge prompt.
Appendix E Detailed Introductions of Safety Benchmarks
AdvBench Zou et al. (2023) contains 521 lines of harmful behaviors, which has been widely adopted for evaluating the vulnerabilities of LLMs and VLMs. We follow previous works Zheng et al. (2024); Zou et al. (2023) to select 100 random prompts for evaluation.
MM-SafetyBench Liu et al. (2024c) is a widely used safety benchmark dataset that consists of $1,680$ questions across 13 safety topics (unsafe scenarios) listed by OpenAI, such as privacy violation, fraudulent, and illegal activities.
RedTeam-2K Luo et al. (2024) is a meticulously curated collection of $2,000$ harmful queries aimed at identifying alignment vulnerabilities in VLMs. It spans 16 safety policies and incorporates queries from 8 distinct sources.
Harmful_Instructions Qi et al. (2024) consists of 40 malicious prompts for evaluating jailbreak attacks.
Appendix F Jailbreak Methods
We attack the target VLM using six types of attacks:
Visual Adversarial Attacks (VAA) Qi et al. (2024) generates visual adversarial examples optimized using a ’few-shot’ corpus to jailbreak VLMs. We follow Qi et al. (2024) and evaluate on the Harmful_Instruction dataset, which contains 40 toxic prompts.
image Jailbreaking Prompt (imgJP) Niu et al. (2024) leverages harmful textual prompts while optimizing the input image to maximize the likelihood of generating a positive affirmation. Following Niu et al. (2024), we select 100 random prompts from AdvBench Zou et al. (2023) for evaluation.
Greedy Coordinate Gradient (GCG) Zou et al. (2023) aims to find a universal adversarial suffix that, when appended to the textual prompt, can jailbreak VLMs. GCG attack is evaluated on 100 random prompts from AdvBench Zou et al. (2023).
AutoDAN Liu et al. (2024b) automatically generates stealthy textual jailbreak prompts using a hierarchical genetic algorithm. The dataset used to evaluate the AutoDAN attack is the same as the dataset evaluating the GCG attack.
Vanilla Attack that directly inputs the jailbreak texts with the clean images into the model. We evaluate our BlueSuffix on defending the vanilla attack using the RedTeam-2k dataset to test the transferability of our defense method across datasets and models.
BAP Attack Ying et al. (2024), a state-of-the-art bimodal attack, which converts the clean images into jailbreak images via universal adversarial perturbation (UAP) while enhances the original jailbreak texts using ChatGPT. We first evaluate our method in defending against the BAP attack on the MM-SafetyBench dataset Liu et al. (2024c). To test the transferability of our defense method across different datasets and models, we also evaluate it on the BAP attack using the RedTeam-2k dataset Luo et al. (2024).
Appendix G Detail Introduction of Baseline Defenses
DiffPure is a diffusion model-based denoising defense that purifies adversarially perturbed images into benign ones. The denoiser is fine-tuned based on a pre-trained diffusion model—specifically, the CLIP image encoder—using pairs of clean and adversarial images. By leveraging the power of generative modeling, DiffPure eliminates various types of adversarial noise. It operates as a preprocessing step independent of the architecture or parameters of the downstream VLM, making it compatible with a wide range of models. Additionally, by focusing on the input distribution rather than specific attack strategies, DiffPure offers robust generalization against diverse adversarial attacks.
Safety Prompt is a defense method designed to optimize the system’s safety prompt using Directed Representation Optimization (DRO). By inserting an optimized safety prompt, the representations of textual inputs tend to move along the ”refusal direction,” thereby safeguarding the target VLM against harmful content.
Robust Refusal Dynamic Defense (R2D2) is adversarial training method for robust refusal, which fine-tunes target models on a dynamic pool of test cases continually updated by a strong optimization-based red teaming method.
Continuous Adversarial Training (CAT) is a fast adversarial training algorithm composed of two losses: the first makes the model robust on continuous embedding attacks computed on an adversarial behavior dataset; the second ensures the usefulness of the final model by fine-tuning on utility data.
VLGuard directly fine-tunes the target VLMs on a safety instruction-following dataset.
Appendix H Perspective API Evaluation
We also employed the Perspective API to score the toxicity of LLaVA model responses under vanilla attacks, BAP attacks, and when defended by BlueSuffix against BAP attacks on the MM-SafetyBench dataset. The Perspective API provides scores ranging from 0 to 1, with higher scores indicating a greater likelihood that the comment is considered toxic. Although the Perspective API’s toxicity scoring is relatively lenient and may not be suitable as a judge for jailbreak detection, it still serves to demonstrate the effectiveness of BlueSuffix as a defense.
The results are as follows:
  • Vanilla Attack: Score of 0.0746
  • BAP Attack: Score of 0.2014
  • BlueSuffix Defense: Score of 0.0271
As a reference, the sample provided by Google, “Friendly greetings from Python”, scored 0.0260, while the score achieved by our defense is 0.0271.
Appendix I Suffix Generator
@@
Figure 7 illustrates six example inputs (three jailbreaks, three benign) purified by our BlueSuffix. Figures 8 and 9 demonstrate that our suffix generator produces diverse outputs while maintaining a high degree of compatibility with benign prompts. In Figure 8, the attacker inputs image-text prompts. Our BlueSuffix purifies the visual prompts and rewrites the textual prompts by appending a suffix. With the purified prompt, the target VLM successfully identifies the malicious content and generates a benign response, including an explanation for refusing to answer. Figure 9 shows a user asking the target VLM benign queries. Our BlueSuffix processes these queries effectively, particularly for the textual prompts. The suffix generator produces a positive response that is relevant to the textual prompt, guiding the target VLM in answering the question. Our method focuses on enabling VLMs to independently evaluate queries, making it more practical for real-world scenarios.
@@
Appendix J An example of Stronger Jailbreak Prompt
Figure 10 illustrates an example of a strong jailbreak textual prompt rewritten by GPT-4o.